You took a 5 year, $100,000 loan. The loan has equal principal payments. The loan carries a 6% annual interest rate and is paid back in annual payments.

1. What is the outstanding balance of the loan after 3 years?

2. Compute an amortization table for the loan.

3. What is the interest payment on the fourth installment?
Answer #1

Working note:

Using financial calculator BA II Plus - Input details:

#

I/Y = Rate/Frequency =

6.000000

FV = Future value =

$0

N = Total payment term x Frequency =

                            5

PV = Present value of Loan =

-$100,000.00

CPT > PMT = Payment =

$23,739.64

Alternate formula-based method:

PMT = Payment = |PV| x R% x (1+R%)^N / ((1+R%)^N - 1)

$23,739.64

Amortization table:

Year

Beginning Balance

Payment

Interest

Repayment of principal

Ending balance

Y

OP

PMT

I = OP x Rate

AM = PMT - I

CB = OP - AM

1

$100,000.00

$23,739.64

                  6,000.00

$17,739.64

$82,260.36

2

                82,260.36

$23,739.64

                  4,935.62

$18,804.02

                63,456.34

3

                63,456.34

$23,739.64

                  3,807.38

$19,932.26

              43,524.08

4

                43,524.08

$23,739.64

                  2,611.44

                21,128.20

                22,395.89

5

                22,395.89

$23,739.64

                  1,343.75

                22,395.89

                              -  

1. What is the outstanding balance of the loan after 3 years?

Answer: Please check the amortization table for 3 years Outstanding Balance which is = $43,524.08

2. Compute an amortization table for the loan.

Year

Beginning Balance

Payment

Interest

Repayment of principal

Ending balance

Y

OP

PMT

I = OP x Rate

AM = PMT - I

CB

1

$100,000.00

$23,739.64

                  6,000.00

$17,739.64

$82,260.36

2

                82,260.36

$23,739.64

                  4,935.62

$18,804.02

                63,456.34

3

                63,456.34

$23,739.64

                  3,807.38

$19,932.26

              43,524.08

4

                43,524.08

$23,739.64

                  2,611.44

                21,128.20

                22,395.89

5

                22,395.89

$23,739.64

                  1,343.75

                22,395.89

                              -  

3. What is the interest payment on the fourth installment?

Answer: Please check amortization table Interest column which is = $2,611.44